# Physics-Informed Weakly Supervised Learning for Interatomic Potentials

### v1

@@ D.4 C–H Potential Energy Profile of Aspirin
Table A11: Performance of PaiNN employed in Fig. 2 (c, d). All the models other than the reference model ($N_{\rm train}=1000$) use the OC20’s hyper-parameters. For the reference model, we tuned the hyper-paramters of PaiNN model following the original paper [34].
@@
This section describes the detailed setup and procedure for Section 5.3. First, we trained PaiNN with and without PIWSL losses using the aspirin data from rMD17 with training set sizes of 100 and 200. For PIWSL, we used $(C_{\rm PITC},C_{\rm PISC},\epsilon_{\rm max})=(1.2,0.01,0.015)$. The other experimental setups are the same as the other rMD17 experiments given in Section B.1. We used the PaiNN model with gradient-based forces to obtain the reference model and tuned the model hyper-parameter with Optuna [56]. The obtained models’ performance is provided in Table A11. Then, we prepared the aspirin molecule structures, including the corresponding atomic coordinates and atomic types. For these structures, we perturbed one of the C-H bonds with a bond length from 0.8 Å to 1.8 Å. We prepared 100 structures and estimated the corresponding potential energy with the pre-trained models. The aspirin data is provided in our publicly available source code.
D.5 Metric Dependence of PITC
@@
Table A12: Metric dependence of PITC. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations. MAE refers to the mean absolute error, and MSE denotes the mean square error.
@@
D.6 Perturbation Magnitude Dependence of PITC
Table A13: Perturbation magnitude dependence of PITC. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ D.6 Perturbation Magnitude Dependence of PITC
In this section, we provide the result of the perturbation magnitude dependence of PIWSL, i.e., $\|\delta\mathbf{r}\|=\epsilon$. For simplicity, we only consider the PITC loss (the coefficient of the PITC and PISC losses are set as $0.1$ and 00). The results are provided in Table A13 and demonstrate that the longer perturbation vector length is fruitful for force predictions. However, values that are too large are harmful to predicting potential energy. In this study, we used the ANI-1x data set and the 1000 training samples. These samples differ from the one used to train the model in the main text to avoid overfitting the test data set.
D.7 Dependence of PITC on the Number of Perturbed Atoms
Table A14: Dependence of PITC on the number of perturbed atoms. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ D.7 Dependence of PITC on the Number of Perturbed Atoms
D.8 Dependence on the Number of Training Iterations
Table A15: Dependence on the number of training iterations. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ D.8 Dependence on the Number of Training Iterations
D.9 Additional Experiments with Gradient-Based Forces
Table A16: Results of PIWSL with gradient-based force predictions. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ D.9 Additional Experiments with Gradient-Based Forces
In this section, we provide the result of the training with the gradient-based force predictions. The results are provided in Table A16 and demonstrate that our PIWSL loss enables a better force prediction, even in the case of gradient-based force predictions. These results also indicate that our PIWSL method can improve the ML model performance in the case of MLIPs commonly applied in computational chemistry and materials science. We consider that this is partly due to the effectiveness of the weak label at $\mathbf{r}+\delta\mathbf{r}$ as indicated by the WL results, which show an improvement of the performance different from the case with the direct force branch (see also subsection 5.5). We hypothesize that the further improvement results from the additional gradient calculation as indicated in Eq. (A5) and Eq. (A6). This observation also indicates that our PIWSL method can potentially improve other generic property prediction tasks by calculating their first derivatives in terms of the atomic coordinate and utilizing the proposed loss functions. In this study, the coefficient of the PITC and PISC losses are set as $0.1$ and $0.01$ with $\epsilon_{\rm max}=0.01$. The weak label loss coefficient is set as $0.1$.
D.10 Experiments with the Weak Label Loss and Predicted instead of Reference Forces
Table A17: Additional results for the weak label loss. WL denotes the weak label loss and WL+FP refers to the weak label loss including force predictions. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
Model Case Baseline WL WL + FP
PaiNN (w RF) E 56.62 $\pm$ 2.80 81.86 $\pm$ 9.39 95.66 $\pm$ 5.90
F 12.96 $\pm$ 0.06 14.54 $\pm$ 0.12 14.55 $\pm$ 0.17
(w/o RF) E 42.36 $\pm$ 0.30 35.79 $\pm$ 0.70 41.77 $\pm$ 4.82
F 24.25 $\pm$ 0.00 24.25 $\pm$ 0.01 24.68 $\pm$ 0.54
Equiformer (w RF) E 54.52 $\pm$ 4.52 31.02 $\pm$ 3.99 29.91 $\pm$ 1.19
F 10.10 $\pm$ 0.00 13.43 $\pm$ 0.92 12.16 $\pm$ 0.31
(w/o RF) E 43.14 $\pm$ 0.86 30.08 $\pm$ 1.97 88.59 $\pm$ 11.36
F 24.25 $\pm$ 0.00 24.25 $\pm$ 0.00 293.41 $\pm$ 26.96
In this section, we provide the additional result of the training with the weak label loss using force predictions, that is, by substituting the MLPI model’s force prediction into $\mathbf{F}^{\rm ref}_{i,\mathcal{S}}$ in Eq. (A3). This is partly due to the inconsistent setup of the experiment in the Table 3 in Section 5.4, in which the weak label loss is calculated using the reference force even for the case without reference force labels. The results are provided in Table A17, which shows that the weak label calculated using reference energy and predicted force does not improve the performance in almost all cases in comparison to the weak label loss. This indicates that the weak label estimated using force prediction only is not enough to estimate sufficiently accurate potential energy at $\mathbf{r}+\delta\mathbf{r}$, and the PITC loss’s relational consistency between the potential energy and force predictions is essential. In this study, the coefficient of the WL loss is set as $1$ and $0.1$ in the case with reference force and without reference force, respectively.
D.11 Reducing Curl of Forces for Models with the Force Branch
Table A18: Curl of forces for models with the force branch. The presented numerical values are the absolute values of the total curl of the force evaluated for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ D.11 Reducing Curl of Forces for Models with the Force Branch
In this section, we study the effect of our loss functions on the curl of forces in the case of the model with the force branch. The results are provided in Table A18, which shows that our PITC loss reduces the curl of the predicted forces, allowing better energy conservation during MD simulations. In this study, we used the ANI-1x data set and the 1000 training samples. These samples differ from the one used to train the model in the main text to avoid overfitting the test data set. The hyper-parameters of the PITC and PISC losses are $(C_{\rm PITC},C_{\rm PISC},\epsilon_{\rm max})=(1.2,0.01,0.025)$.
Appendix E Potential Broader Impact and Ethical Aspects

### v2

@@ E.4.1 C–H Potential Energy Profile of Aspirin
Table A13: Performance of PaiNN employed in figure 2 (c, d). All the models other than the reference model ($N_{\rm train}=1000$) use the OC20’s hyper-parameters. For the reference model, we tuned the hyper-paramters of PaiNN model following the original paper (Schütt et al., 2021).
@@
This section describes the detailed setup and procedure for section 5.3. First, we trained PaiNN with and without PIWSL losses using the aspirin data from rMD17 with training set sizes of 100 and 200. For PIWSL, we used $(C_{\rm PITC},C_{\rm PISC},\epsilon_{\rm max})=(1.2,0.01,0.015)$. The other experimental setups are the same as for rMD17 experiments presented in section C.1. We used the PaiNN model with gradient-based forces to obtain the reference model and tuned the model hyper-parameter with Optuna (Akiba et al., 2019). The obtained models’ performance is provided in Table A13. Then, we prepared the aspirin molecule structures, including the corresponding atomic coordinates and atomic types. For these structures, we perturbed one of the C-H bonds with a bond length from 0.8 Å to 1.8 Å. We prepared 100 structures and estimated the corresponding potential energy with the pre-trained models. The aspirin data is provided in our publicly available source code.
E.4.2 MD Simulation Stability Analysis
NVE-Ensembles
This section describes the detailed setup and the procedure for our analysis of MD simulations in section 5.3. Because our implementation builds upon the source code provided by Fu et al. (2023), we used their scripts for performing MD simulations. However, we added a minor modification to enable MD simulations in the microcanonical (NVE) statistical ensemble, i.e., the particle position and velocity are updated with velocity Verlet algorithm(Verlet, 1967)  [^7] Note that the total energy conservation necessary for the microcanonical statistical ensemble is in general not perfectly satisfied due to the numerical error, in particular, when the force is not calculated as the curl of the force.. We set the initial temperature to 300 K and the integration time step to 0.5 fs for all simulations. As defined by Fu et al. (2023), the stability of an MD simulation for a target molecule is defined as the time $T$ during which the bond lengths satisfy the following condition
$\max_{(i,j)\in\mathcal{B}}|(||{\bf x}_{i}(T)-{\bf x}_{j}(T)||)-b_{i,j}|>\Delta\,,$ (A17)
where $\mathcal{B}$ denotes the set of all bonds, $\{i,j\}$ denote the pair of bonded atoms, and $b_{i,j}$ denotes the equilibrium bond length. Following Fu et al. (2023), we set $\Delta=0.5\text{\AA}$. This definition indicates when the molecule experiences significant structural changes during the MD simulation.
We trained PaiNN and Equiformer v2 models with and without PIWSL losses using the aspirin data from rMD17. We used training set sizes of 100 and 200 for PaiNN-GF and 1000 for PaiNN and Equiformer v2 with direct force. The corresponding stability values are presented in Table A14. The hyperparameters for the PIWSL loss are provided in Table A15. To train direct force model with 1000 samples, we used second-order PITC and PISC losses because of their good accuracy. The results for the stability of the PaiNN (direct and gradient-based force) and Equiformer v2 models are shown in Table A16. To select the hyperparameters in Table A15, a series of MD simulations with a fixed random number seed for the initial atomic velocities was used. For our final stability results in Table A14, a series of MD simulations with three different random seeds for velocities and the selected hyperparameters was used. figure A1 presents the temporal evolution of the total energy during the MD simulations reported in figure 3, which shows that the total energy is conserved within 0.05% deviation until the disruption. These results indicate that PIWSL slightly enhances the energy conservation capability of MLIP models in most cases, though the improvement is relatively modest due to the small deviations observed.
NVT-Ensembles
To investigate the effect of thermostats, we also performed MD simulations in the canonical (NVT) statistical ensemble, where temperature is maintained constant. To keep the temperature constant, we used Nosé-Hoover thermostat (Nosé, 1984; Hoover, 1985). The initial and target temperatures were both set to 300 K for all simulations. The integration time step was set to 0.5 fs and the characteristic parameter $\tau$ for the thermostat is set to 20 fs. The result, shown in figure A2, demonstrate that the thermostat stabilizes the simulations by mitigating the increase in kinetic energy.
Table A14: Stability of the models employed in the MD analysis. The presented numerical values are the stability defined by Eq. (A17) measured in ps. The results are obtained as an average over three different random seeds for the initial velocity of the atoms in the target aspirin molecule. "GF" denotes the gradient-based force prediction.
$N_{\mathrm{train}}=100$ $N_{\mathrm{train}}=200$ $N_{\mathrm{train}}=1000$
PaiNN Baseline 2.68 $\pm$ 0.13
PIWSL 4.65 $\pm$ 0.72
Equiformer Baseline 14.53 $\pm$ 8.61
PIWSL 24.65 $\pm$ 8.24
PaiNN-GF Baseline 3.25 $\pm$ 3.98 220.5 $\pm$ 137.7
PIWSL 15.07 $\pm$ 10.09 267.7 $\pm$ 56.0
Table A15: Hyper-parameters for the PIWSL loss used in the MD analysis. We used the following hyper-parameter for MD simulation analysis: ($C_{\rm PITC},C_{\rm PISC},\epsilon_{\mathrm{max}})$= Case $\alpha$: (0.01, 0.001, 0.025), Case $\beta$: (1.2, 0.01, 0.01), Case $\gamma$: (1.2, 0.01, 0.025), Case $\delta$: (1.2, 0.01, 0.015), Case $\epsilon$: (0.1, 0.01, 0.01), and Case $\zeta$: (1.0, 0., 0.01). "GF" denotes the gradient-based force prediction.
Dataset Size Equiformer v2 PaiNN PaiNN-GF
rMD17 100 $\alpha$ $\delta$ $\epsilon$
(Aspirin) 200 $\beta$ $\beta$ $\zeta$
1000 $\epsilon$ $\gamma$
Table A16: Energy and force errors for the models employed in the MD analysis. The presented numerical values are the root-mean-square errors (RMSEs) of energy (E) and force (F). Energy RMSE is given in kcal/mol, while force RMSE is in kcal/mol/Å. "GF" denotes the gradient-based force prediction.
$N_{\mathrm{train}}=100$ $N_{\mathrm{train}}=200$ $N_{\mathrm{train}}=1000$
Baseline PIWSL Baseline PIWSL Baseline PIWSL
PaiNN-DF E 6.55 5.64 5.11 4.48 2.30 0.99
F 7.38 7.36 3.95 3.97 1.63 1.61
Equiformer v2 E 4.79 4.64 4.92 4.82 1.39 1.19
F 4.86 4.90 2.50 2.42 0.76 0.75
PaiNN-GF E 6.05 6.03 6.01 6.02
F 6.41 6.33 3.50 3.53
Figure A1: Analysis of the total energy conservation using MD simulations with MLIP models. The amount of the change of the total energy during MD simulations is assessed for the baseline MLIP models and those trained with PIWSL. The total energy is measured at the initial and final time-step and the difference is normalized by the total energy at the initial time-step. All results are obtained for the aspirin molecule and MD simulations in the microcanonical (N V E) statistical ensemble.
Figure A2: Stability analysis of the MLIP models during MD simulations. Stability during MD simulations is assessed for the baseline MLIP models and those trained with PIWSL. All results are obtained for the aspirin molecule and MD simulations in the canonical (N V T) statistical ensemble. We measure stability during MD simulations according to (Fu et al., 2023).
E.5 Training setup for MD17-CCSD(T) Experiments
In this section, we provide the training setup of finetuning of MACE-OFF discussed in section 5.4. The considered model is MACE-OFF (large) (Kovács et al., 2023) which is the MACE model pretrained on SPICE (Eastman et al., 2023), QMugs (Isert et al., 2022), and liquid water (Schran et al., 2021) datasets. The foundation model is finetuned on the aspirin molecule data in CCSD dataset (Chmiela et al., 2018) whose potential energy is obtained at quantum-chemical CCSD level of accuracy. The data includes 1500 samples which are splitted into 950/50/500 as train/validation/test datasets. To emulate general coupling-cluster method dataset, we only use potential energy label for training. We utilize the set up provided in the official repository [^8] https://mace-docs.readthedocs.io/en/latest/guide/finetuning.html with modifying learning rate from $10^{-2}$ to $10^{-3}$, to improve the performance. For the model trained scratch, we use the same architecture with setting energy weight to 40 and learning rate to $10^{-3}$ to improve the performance. For PIWSL, we used the second-order PITC and 2pt-with-label PISC losses Equation A10 with the following parameters: ($C_{\rm PITC},C_{\rm PISC},\epsilon)=(0.55,4.5,0.01)$.
E.6 Metric Dependence of PITC
@@
Table A17: Metric dependence of PITC. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations. MAE refers to the mean absolute error, and MSE denotes the mean square error.
@@
E.7 Perturbation Magnitude Dependence of PITC
Table A18: Perturbation magnitude dependence of PITC. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ E.7 Perturbation Magnitude Dependence of PITC
In this section, we provide the result of the perturbation magnitude dependence of PIWSL, i.e., $\|\delta\mathbf{r}\|=\epsilon$. For simplicity, we only consider the PITC loss (the coefficient of the PITC and PISC losses are set as $0.1$ and $0.0$). The results are provided in Table A18 and demonstrate that the longer perturbation vector length is fruitful for force predictions. However, values that are too large are harmful to predicting potential energy. In this study, we used the ANI-1x data set and the 1000 training samples. These samples differ from the one used to train the model in the main text to avoid overfitting the test data set.
E.8 Dependence of PITC on the Number of Perturbed Atoms
Table A19: Dependence of PITC on the number of perturbed atoms. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ E.8 Dependence of PITC on the Number of Perturbed Atoms
E.9 Dependence on the Number of Training Iterations
Table A20: Dependence on the number of training iterations. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ E.9 Dependence on the Number of Training Iterations
E.10 Additional Experiments with Gradient-Based Forces
Table A21: Results of PIWSL with gradient-based force predictions. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ E.10 Additional Experiments with Gradient-Based Forces
In this section, we provide the result of the training with the gradient-based force predictions. The results are provided in Table A21 and demonstrate that our PIWSL loss enables a better force prediction, even in the case of gradient-based force predictions. These results also indicate that our PIWSL method can improve the ML model performance in the case of MLIPs commonly applied in computational chemistry and materials science. We consider that this is partly due to the effectiveness of the weak label at $\mathbf{r}+\delta\mathbf{r}$ as indicated by the WL results, which show an improvement of the performance different from the case with the direct force branch (see also section F). We hypothesize that the further improvement results from the additional gradient calculation as indicated in Eq. (A5) and Eq. (A6). This observation also indicates that our PIWSL method can potentially improve other generic property prediction tasks by calculating their first derivatives in terms of the atomic coordinate and utilizing the proposed loss functions. In this study, the coefficient of the PITC and PISC losses are set as $0.1$ and $0.01$ with $\epsilon_{\rm max}=0.01$. The weak label loss coefficient is set as $0.1$.
E.11 Reducing Curl of Forces for Models with the Force Branch
Table A22: Curl of forces for models with the force branch. The presented numerical values are the absolute values of the total curl of the force evaluated for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations.
@@ E.11 Reducing Curl of Forces for Models with the Force Branch
In this section, we study the effect of our loss functions on the curl of forces in the case of the model with the force branch. The results are provided in Table A22, which shows that our PITC loss reduces the curl of the predicted forces, allowing potentially better energy conservation during MD simulations. In this study, we used the ANI-1x data set and the 1000 training samples. These samples differ from the one used to train the model in the main text to avoid overfitting the test data set. The hyper-parameters of the PITC and PISC losses are $(C_{\rm PITC},C_{\rm PISC},\epsilon_{\rm max})=(1.2,0.01,0.025)$. It is theoretically possible to define a loss function aimed at directly minimizing the absolute value of the curl of forces. However, this approach necessitates calculating the Hessian matrix, which requires a substantial memory cost given the limitations of current computational resources. Developing a method to train with such a loss function while mitigating memory requirements is a promising direction for future research.
Appendix F Further Analyses of PIWSL
F.1 Training MLIPs without Reference Forces
In the following, we explore scenarios where only potential energy labels are available. This situation commonly arises when calculating energy labels with chemically accurate approaches, such as CCSD(T)/CBS (Hobza & Šponer, 2002; Feller et al., 2006), for which force calculation is infeasible. To consider practical applications, we examine two cases: (1) predicting force by a force branch (FB) and (2) predicting force as a gradient of the potential energy (GF). The former enables fast force prediction and is popular in the machine learning community, while the latter requires additional gradient calculation but yields curl-free force predictions. It is popular in computational chemistry as it ensures the conservation of the total energy during MD simulations. The results are provided in Table A23; training without reference forces is achieved by setting the relative force contribution to zero in Eq. (3). The PIWSL method consistently performs better than the baseline for the FB and GF cases. However, a more significant improvement in the force prediction performance is observed in the GF case. We attribute this phenomenon to the inherent nature of PIWSL, which requires consistency between the potential energy and atomic forces, as discussed in section 5.3. This result aligns with our expectations, confirming the capability of our PIWSL method to enable ML models to reduce the error in the predicted forces. Overall, PIWSL opens a new possibility for training MLIP models using highly accurate reference methods, such as CCSD(T)/CBS.
Table A23: Results for models trained on the ANI-1x data set without reference forces. All models are trained using 1000 training samples. FB refers to the setting where the force branch estimates the force, and GF denotes the setting where the force is estimated by the gradient of the potential energy with respect to the atomic coordinates.
Model Case Baseline PIWSL
PaiNN FB E 42.36 $\pm$ 0.30 25.42 $\pm$ 0.72
F 24.25 $\pm$ 0.00 20.54 $\pm$ 0.08
GF E 41.83 $\pm$ 1.81 29.71 $\pm$ 0.55
F 83.36 $\pm$ 2.85 24.02 $\pm$ 0.95
Equiformer FB E 43.14 $\pm$ 0.86 29.48 $\pm$ 0.51
F 24.25 $\pm$ 0.00 21.99 $\pm$ 0.49
GF E 42.55 $\pm$ 0.99 32.66 $\pm$ 1.11
F 35.70 $\pm$ 0.78 21.83 $\pm$ 0.27
F.2 Comparing PITC with the Taylor-Expansion-Based Weak Label Loss.
We compare the PIWSL method with the Taylor-expansion-based weak label (WL) approach (Cooper et al., 2020), whose loss function is presented in Eq. (A3). For simplicity, we only consider the PITC loss in Eq. (4.1). For a fair comparison, we consider the following two cases. First, we train with reference forces and energies (w. RF). Second, we train the methods without reference forces and use only the reference energies. For the training with reference forces, we set the numeric coefficient of the PITC loss to $1.0$; for the training without reference forces, the coefficient is set to $0.1$. Note that the WL loss without the reference force is calculated using the predicted force labels. The results are provided in Table A24.
Our PITC loss demonstrates the best accuracy in all cases, with and without the reference forces. Interestingly, PaiNN failed to learn the potential energy with the WL loss and reference forces. We hypothesize this to be due to the imbalance of the training between the energies and forces. Specifically, the WL loss trains only the potential energy, resulting in an inconsistency between the energy and force branches, which share the same readout layer that experiences more frequent updates using the potential energy. This hypothesis is supported by the results for the training without reference forces, where the error in energy is reduced compared to the baseline. A further validation in a similar experiment in the case of GF is provided in section E.10. However, the proposed PITC loss still performs better here. In summary, the PITC loss enables MLIPs to learn energies and forces consistent with each other and does it better than the previously proposed WL method.
Table A24: Comparison of PITC and the Taylor-expansion-based weak label loss. WL (+FP) denotes the Taylor-expansion-based method using reference energies and either reference (w. RF) or predicted (w/o. RF) forces; see Eq. (A3). The listed values are the RMSE values for energies in kcal/mol and atomic forces in kcal/mol/Å. All models are trained on the ANI-1x data set using 1000 configurations, with (w.) and without (w/o.) reference atomic forces (RF).
Model Case Baseline PITC WL (+FP)
PaiNN (w. RF) E 56.62 $\pm$ 2.80 30.94$\pm$ 0.56 81.86$\pm$ 9.39
F 12.96$\pm$ 0.06 12.04$\pm$ 0.04 14.54$\pm$ 0.12
(w/o. RF) E 42.36$\pm$ 0.30 25.42$\pm$ 0.72 41.77$\pm$ 4.82
F 24.25$\pm$ 0.00 20.54$\pm$ 0.08 24.68$\pm$ 0.54
Equiformer (w. RF) E 54.52$\pm$ 4.52 23.16$\pm$ 0.19 31.02$\pm$ 3.99
F 10.10$\pm$ 0.00 10.03$\pm$ 0.05 13.43$\pm$ 0.92
(w/o. RF) E 43.14$\pm$ 0.86 29.48$\pm$ 0.51 88.59$\pm$ 11.36
F 24.25$\pm$ 0.00 21.99$\pm$ 0.49 293.41$\pm$ 26.96
F.3 Ablating the Impact of PITC and PISC Losses.
We conduct an ablation experiment to analyze the impact of PITC and PISC losses. Results in Table F.3 indicate that the PITC loss predominantly improves the accuracy of resulting models, especially for PaiNN. Using just the PISC loss does not consistently improve accuracy but stabilizes training when combined with PITC. This combined approach notably benefits Equiformer v2. For Equiformer v2, we repeated the experiment five times to reduce the effect from an outlier on the PITC loss.
Table A25: Results for models with direct-force trained on the ANI-1x data set with ablated weakly supervised losses. All models are trained using 1000 training samples. All results are obtained by averaging over three independent runs. Energy RMSE is given in kcal/mol, while force RMSE is given in kcal/mol/Å.
Model PITC PISC E F
PaiNN 56.62 $\pm$ 2.80 12.96 $\pm$ 0.06
24.60 $\pm$ 0.18 11.51 $\pm$ 0.03
58.30 $\pm$ 2.10 13.18 $\pm$ 0.29
24.53 $\pm$ 0.48 11.43 $\pm$ 0.05
Equiformer 54.52 $\pm$ 4.52 10.10 $\pm$ 0.00
32.64 $\pm$ 26.48 9.64 $\pm$ 0.03
48.96 $\pm$ 4.96 10.30 $\pm$ 0.06
20.89 $\pm$ 0.50 9.68 $\pm$ 0.03
F.4 Adversarial Directions for Perturbing Atomic Positions.
The following discusses the dependence of the PIWSL’s performance on selecting the vector $\delta\mathbf{r}$ in Eq. (4.3) employed to perturb atomic positions. The detailed implementation and setups are provided in section C.1. Table A26 compares the results obtained for a randomly-sampled vector $\delta\mathbf{r}$ and for the one determined adversarially. The results demonstrate that both approaches improve the performance compared to the baseline without weak supervision, though the results might depend on the employed model.
Table A26: PIWSL’s performance dependence on the atomic position perturbation vector. The numerical values are RMSEs for the energy in kcal/mol and force in kcal/mol/Å. All results are provided for the ANI-1x data set and models trained using 1000 configurations.
Baseline Random (Eq. (6)) Adversarial (Eq. (7))
PaiNN E 56.62 $\pm$ 2.80 24.53 $\pm$ 0.48 33.67 $\pm$ 1.12
F 12.96 $\pm$ 0.18 11.43 $\pm$ 0.05 12.74 $\pm$ 0.14
Equiformer E 54.52 $\pm$ 4.52 23.16 $\pm$ 0.50 20.54 $\pm$ 0.21
F 10.10 $\pm$ 0.00 10.03 $\pm$ 0.03 9.93 $\pm$ 0.04
F.5 Performance Difference Between First and Second order PIWSL
In this section, we present a comparison between first- and second-order PIWSL methods. We evaluate a foundation model, MACE-OFF, fine-tuned on the MD17 Aspirin dataset with CCSC(T) labels and the MD22 buckyball catcher data set, as well as PaiNN-GF, a gradient-based force model trained on 1,000 samples from ANI-1x. The results are provided in Table A27.
Our findings indicate that the second-order term becomes more significant as model performance improves, particularly when force prediction error becomes small. This suggests that the first-order term, which encourages the model to predict the potential energy of neighboring conformations based on a first-order Taylor expansion, plays a crucial role in promoting a smoother potential energy surface. This effect arises because, in MACE and PaiNN-GF models, force predictions are derived as the gradient of the potential energy surface.
Table A27: Performance difference between first and second order PIWSL. The numerical values are RMSEs (MD17-CC) and MAE (MD22-BB) for the energy in kcal/mol and force in kcal/mol/Å. MD17-CC denotes MD17 aspirin data with CCSD label, MD22-BB denotes MD22 buckyball catcher data, and ANI-1x (1K) denotes 1000 samples from ANI-1x.
Dataset Model Baseline 1st-order PIWSL 2nd-order PIWSL
MD17-CC MACE-OFF E 1.21 $\pm$ 0.00 0.90 $\pm$ 0.03 0.72 $\pm$ 0.01
F 6.90 $\pm$ 0.01 4.59 $\pm$ 0.16 3.77 $\pm$ 0.13
MD22-BB MACE-OFF E 1.16 $\pm$ 0.15 1.16 $\pm$ 0.28 0.99 $\pm$ 0.05
F 0.35 $\pm$ 0.00 0.35 $\pm$ 0.00 0.34 $\pm$ 0.00
ANI-1x (1K) PaiNN-GF E 23.57 $\pm$ 0.62 18.62 $\pm$ 0.09 18.88 $\pm$ 0.23
F 11.32 $\pm$ 0.08 10.94 $\pm$ 0.01 11.01 $\pm$ 0.00